OpenStudy (anonymous):

How many combinations are there for a 3-digit number?

OpenStudy (anonymous):

I think it's 10^3 , while my sister thinks it's 10P3...

OpenStudy (anonymous):

Oh, the hundred place can be 0 btw

OpenStudy (anonymous):

That means 012 is acceptable.

hartnn (hartnn):

then its 10^3 from 0 to 999 = 999-0+1 = 1000

OpenStudy (anonymous):

Nice! Thanks :D abc 10 choices x 10 choices x 10 choices = 10^3, right?!

hartnn (hartnn):

absolutely...

OpenStudy (anonymous):

Thanks again!! :)

hartnn (hartnn):

welcome again ^_^